# IPPP-25-XXX

### v1

@@ Abstract
I review the standard instanton framework for determining the $\theta$-dependence of instanton-dominated correlation functions in QCD. I then contrast these well-established semiclassical results with the recent assertion of Refs. [Ai:2020ptm, Ai:2024cnp], that $\theta$-phases are absent and that strong interactions preserve CP.
@@
Consider QCD with massive quarks and include the conventionally normalised $\theta$-term, $\frac{i\theta}{16\pi^{2}}\,{\rm tr}\,F_{\mu\nu}\tilde{F}_{\mu\nu}$, in the Euclidean Lagrangian. Assume that the theory is weakly coupled so that a semiclassical treatment is reliable. Our goal is to compute the partition function $Z(\theta)$ and thereby determine the vacuum energy density as a function of $\theta$,
@@
where $V_{4}$ is the four-dimensional Euclidean volume. To proceed, we further assume that the dominant contributions to the path integral for $Z(\theta)$ arise from expanding around classical field configurations consisting of $n+\bar{n}$ widely separated, non-interacting instantons and anti-instantons. Summing over all possible numbers $n$ of instantons and $\bar{n}$ of anti-instantons leads to the standard dilute instanton gas approximation (DIGA) formula [Callan:1976je, Coleman1985Aspects],
@@
Let us now examine more closely the structure of a generic connected correlator in an instanton background, starting from the expression on the right-hand side of (10). We have
@@
$\langle{\rm Correlator}\rangle_{n,\bar{n}}$ $=$ $\langle{\rm Correlator}\rangle_{1}\,Z_{n-1,\bar{n}}\,+\,\langle{\rm Correlator}\rangle_{\bar{1}}\,Z_{n,\bar{n}-1}\,+\,\langle{\rm Correlator}\rangle_{0}\,Z_{n,\bar{n}}.$ (13)
We see that each of these expressions corresponds to an $(n+\bar{n})$-fold product of elementary single-instanton and single-anti-instanton contributions. For example, the expression on the right-hand side of (12) is simply the product of $n$ single-instanton factors and $\bar{n}$ single-anti-instanton factors (divided by the appropriate Bose–Einstein symmetry factor). Likewise, the first term on the right-hand side of (13) represents the one-instanton contribution to the correlator multiplied by (n1)(n-1) disconnected single-instanton contributions and by $\bar{n}$ single-anti-instanton contributions to the partition function. All such terms are disconnected diagrams built from $(n+\bar{n})$ instanton ‘blobs’, exactly as expected in the dilute instanton gas picture.
@@
    When interactions are included, additional genuine multi-instanton contributions can appear in (14). These correspond to true multi-instanton ADHM solutions, whose effects have been explicitly computed in $N=2$ and $N=4$ SYM theories in Refs. [Dorey:1996hu, Nekrasov:2002qd, Dorey:1999pd]. Such multi-instanton contributions are unrelated to the dilute gas approximation; rather, they represent systematic $k$-instanton corrections in the sense of the virial expansion (8). These effects do not play a role in our present analysis of $\theta$-dependence to leading order in the instanton number.
Now we are ready to compare the outlined above standard calculation of $\theta$-dependence to an alternative approach advocated in Refs. [Ai:2020ptm, Ai:2024cnp].
@@
The authors of [Ai:2020ptm, Ai:2024cnp] proposed to re-organise the infinite two-fold summation in the DIGA formula (11) into two steps, and apply the infinite volume limit after completing the first sum but before the second.
@@
$\langle{\rm Correlator}\rangle_{\rm conn}\,=\,\frac{\langle{\rm Correlator}\rangle_{\rm DIGA}}{Z_{\rm DIGA}}\,=\,\,\lim_{N\to\infty}\frac{\sum_{\Delta n=-N}^{N}\sum_{n}\,\langle{\rm Correlator}\rangle_{n,\,\bar{n}=n+\Delta n}}{\sum_{\Delta n=-N}^{N}\sum_{n}\,Z_{n,\bar{n}}}.$ (16)
Performing the inner summations over $n$ at fixed $\Delta n$ gives,
$e^{i\Delta n\bar{\theta}}\left(I_{\Delta n+1}(x)\,\langle{\rm Correlator}\rangle_{1,\,\bar{\theta}=0}\,+\,I_{\Delta n-1}(x)\,\langle{\rm Correlator}\rangle_{\bar{1},\,\bar{\theta}=0}\,+\,I_{\Delta n}(x)\,\langle{\rm Correlator}\rangle_{0}\right),$ (17)
$e^{i\Delta n\bar{\theta}}\,I_{\Delta n}(x),$ (18)
for the expressions in the numerator and the denominator, respectively. Here $I_{\Delta n\pm 1}(x)$ and $I_{\Delta n}(x)$ are the modified Bessel functions and their argument is $x=2V_{4}\,e^{-8\pi^{2}/g^{2}}|K|$.
If we sum these expressions over $\Delta n$ and evaluate the ratio on the right-hand side of (16), we recover the result (14) for the connected correlator $\langle{\rm Correlator}\rangle_{\rm conn}$.
The approach advocated in Refs. [Ai:2020ptm, Ai:2024cnp] is different. These authors insist on taking the infinite volume limit before the infinite sum over $\Delta n$. Specifically, they define the correlator,
$\langle{\rm Correlator}\rangle_{\rm ACGT}\,=\,\lim_{N\to\infty}\,\lim_{V_{4}\to\infty}\frac{\sum_{\Delta n=-N}^{N}\sum_{n}\,\langle{\rm Correlator}\rangle_{n,\,\bar{n}=n+\Delta n}}{\sum_{\Delta n=-N}^{N}\sum_{n}\,Z_{n,\bar{n}}}$ (19)
Mathematically, there is no need to impose the infinite-volume limit either before or after performing the infinite sum: as already noted, the sums in (4) and (11) converge for any finite value of $x$, and moreover the dependence on the spacetime volume cancels in the ratio appearing in (15). Nevertheless, the authors of Refs. [Ai:2020ptm, Ai:2024cnp] were strongly drawn to the idea of imposing the limit $\lim_{V_{4}\to\infty}$, and do it prior to extending the summation over $\Delta n$ to an infinite domain – as prescribed in (19). In these settings, one makes use of the asymptotic values of the Bessel functions,
@@
so that all ratios of $I_{\Delta n\pm 1}(x)/I_{m}(x)$ and $I_{\Delta n}(x)/I_{m}(x)$ are =1=1 in the $V_{4}\to\infty$ limit. Normalising the numerator and denominator expressions (17)-(18) by the same $I_{m}(x)$ and taking the limit $x\to\infty$, one obtains for the numerator of (19):
@@
The authors of Refs. [Ai:2020ptm, Ai:2024cnp] take this cancellation to imply that physical observables, as defined by the order of limits in (19), are insensitive to the CP-violating parameter $\theta$. This assertion forms the core of their claim that CP violation is absent in QCD.
@@
    Since the framework of Refs. [Ai:2020ptm, Ai:2024cnp] does not attempt to go beyond the ideal, non-interacting instanton gas and does not incorporate genuine multi-instanton contributions, it is particularly puzzling that the instanton correlators arising in their construction differ from the leading, single-instanton contributions on the right-hand side of (14).
@@
    The rationale in Refs. [Ai:2020ptm, Ai:2024cnp] for taking the large-$V_{4}$ limit before summing over all topological sectors $\Delta n$ was to localise instantons well inside $V_{4}$, sufficiently far from the boundary at infinity. This was intended to ensure that, in every topological sector, the gauge fields approach the required pure-gauge boundary conditions.
@@
    Beyond QCD and strong CP: Instanton contributions in 4-dimensional gauge theories always carry complex phases which are naturally organised as a Fourier series in $q=e^{2\pi i\tau}$, where $\tau=\frac{\theta}{2\pi}+\frac{4\pi i}{g^{2}}$. In $\mathcal{N}=2$ Seiberg–Witten theory [Seiberg:1994rs] the low-energy effective action is encoded in the holomorphic prepotential $\mathcal{F}(a,\tau)$. For large $a$ one obtains an expansion
    $\mathcal{F}(a,\tau)=\mathcal{F}_{\mathrm{class+1loop}}(a,\tau)+\sum_{k\geq 1}\mathcal{F}_{k}(a)\,e^{2\pi ik\tau}.$ (28)
    where the summation is over all multi-instanton numbers $k$. The complex phases $e^{2\pi ik\tau}$ of instanton contributions [Dorey:1996hu, Nekrasov:2002qd] are essential to ensure holomorphy and correct monodromies of the prepotential [Seiberg:1994rs].
@@
    In $\mathcal{N}=4$ SYM the factors of $e^{2\pi ik\tau}$ allow non-perturbative quantities to be organised into genuine modular or quasi-modular forms [Banks:1998nr, Green:1997tv] and [Dorey:1999pd]. They implement S-duality by providing the correct transformation behaviour under $SL(2,\mathbb{Z})$.
In this note I have focused exclusively on the instanton-based derivation of the $\theta$-dependence, leaving aside other complementary formulations and methods. For a recent ongoing discussion of these broader aspects, see Refs. [Benabou:2025viy, Ai:2025quf].

### v2

# A note on instantons, $\theta$-dependence and strong CP

@@ Abstract
I review the standard instanton framework for determining the $\theta$-dependence of instanton-dominated correlation functions in QCD. I then contrast these well-established semiclassical results with the recent assertion of Refs. [3, 4] that $\theta$-phases are absent and that strong interactions preserve CP for all values of $\theta$. In all scenarios considered, the $\theta$-dependence is shown to be intrinsically non-trivial, being governed either by instantons in the weak-coupling regime or by alternative, non-perturbative considerations when the theory is strongly coupled.
@@
Consider QCD with massive quarks and include the conventionally normalised $\theta$-term, $\frac{i\theta}{16\pi^{2}}\,{\rm tr}\,F_{\mu\nu}\tilde{F}_{\mu\nu}$, in the Euclidean Lagrangian. Assume that the theory is weakly coupled so that a semiclassical treatment is reliable. [Situations in which this assumption fails are discussed at the end of this paper.] Our goal is to compute the partition function $Z(\theta)$ and thereby determine the vacuum energy density as a function of $\theta$,
@@
where $V_{4}$ is the four-dimensional Euclidean volume. To proceed, we further assume that the dominant contributions to the path integral for $Z(\theta)$ arise from expanding around classical field configurations consisting of $n+\bar{n}$ widely separated, non-interacting instantons and anti-instantons. Summing over all possible numbers $n$ of instantons and $\bar{n}$ of anti-instantons leads to the standard dilute instanton gas approximation (DIGA) formula [8, 21, 9],
@@
Let us now examine more closely the structure of a connected correlator in an instanton background. For a direct comparison with the ACGT approach, we consider the two-point correlation function of a massive quark in QCD (as in Eq. (26) of Ref. [4]),
$\langle{\rm Correlator}\rangle_{\rm conn}$ $\equiv$ $\langle\psi(x_{i})\bar{\psi}(x_{j})\rangle_{\rm conn}\,=\,-\frac{\delta^{2}\log Z[\eta,\bar{\eta}]}{\delta\bar{\eta}(x_{i})\delta\eta(x_{j})}|_{\bar{\eta}=0=\eta}\,=\,\frac{\int DA_{\mu}D\bar{\psi}D\psi\,\psi(x_{i})\bar{\psi}(x_{j})\,e^{-S}}{\int DA_{\mu}D\bar{\psi}D\psi\,e^{-S}}$ (11)
and compute the right hand side in the DIGA approximation. We have
@@
$\langle{\rm Correlator}\rangle_{n,\bar{n}}$ $=$ $\langle{\rm Correlator}\rangle_{1}\,Z_{n-1,\bar{n}}\,+\,\langle{\rm Correlator}\rangle_{\bar{1}}\,Z_{n,\bar{n}-1}\,+\,\langle{\rm Correlator}\rangle_{0}\,Z_{n,\bar{n}},$ (14)
where $\langle{\rm Correlator}\rangle_{1}$, $\langle{\rm Correlator}\rangle_{\bar{1}}$ and $\langle{\rm Correlator}\rangle_{0}$ denote $\langle\psi(x_{i})\bar{\psi}(x_{j})\rangle$ computed in the background of a single instanton, single anti-instanton, and in perturbation theory respectively.
We see that each of the expressions in (13) and (14) corresponds to an $(n+\bar{n})$-fold product of elementary single-instanton and single-anti-instanton contributions. For example, the expression on the right-hand side of (13) is simply the product of $n$ single-instanton factors and $\bar{n}$ single-anti-instanton factors (divided by the appropriate Bose–Einstein symmetry factor). Likewise, the first term on the right-hand side of (14) represents the one-instanton contribution to the correlator multiplied by (n1)(n-1) disconnected single-instanton contributions and by $\bar{n}$ single-anti-instanton contributions to the partition function. All such terms are disconnected diagrams built from $(n+\bar{n})$ instanton ‘blobs’, exactly as expected in the dilute instanton gas picture.
@@
    When interactions are included, additional genuine multi-instanton contributions can appear in (15). These correspond to true multi-instanton ADHM solutions, whose effects have been explicitly computed in $N=2$ and $N=4$ SYM theories in Refs. [17, 16, 25, 15]. Such multi-instanton contributions are unrelated to the dilute gas approximation; rather, they represent systematic $k$-instanton corrections in the sense of the virial expansion (8). These effects do not play a role in our present analysis of $\theta$-dependence to leading order in the instanton number.
Now we are ready to compare the outlined above standard calculation of $\theta$-dependence to an alternative approach advocated in Refs. [3, 4].
@@
The authors of [3, 4] proposed to re-organise the infinite two-fold summation in the DIGA formula (12) into two steps, and apply the infinite volume limit after completing the first sum but before the second.
@@
$\langle{\rm Correlator}\rangle_{\rm conn}\,=\,\frac{\langle{\rm Correlator}\rangle_{\rm DIGA}}{Z_{\rm DIGA}}\,=\,\,\lim_{N\to\infty}\frac{\sum_{\Delta n=-N}^{N}\sum_{n}\,\langle{\rm Correlator}\rangle_{n,\,\bar{n}=n-\Delta n}}{\sum_{\Delta n=-N}^{N}\sum_{n}\,Z_{n,\,\bar{n}=n-\Delta n}}.$ (17)
Performing the inner summations over $n$ at fixed $\Delta n$ in the numerator and the denominator in (17) we find,
$\sum_{n}\,$ $\langle{\rm Correlator}\rangle_{n,\,\bar{n}=n-\Delta n}\,=\,$ (18)
$e^{i\Delta n\bar{\theta}}\left(I_{\Delta n-1}(x)\,\langle{\rm Correlator}\rangle_{1,\,\bar{\theta}=0}\,+\,I_{\Delta n+1}(x)\,\langle{\rm Correlator}\rangle_{\bar{1},\,\bar{\theta}=0}\,+\,I_{\Delta n}(x)\,\langle{\rm Correlator}\rangle_{0}\right),$
and
$\sum_{n}\,Z_{n,\,\bar{n}=n-\Delta n}\,=\,e^{i\Delta n\bar{\theta}}\,I_{\Delta n}(x).$ (19)
Here $I_{\Delta n\pm 1}(x)$ and $I_{\Delta n}(x)$ are the modified Bessel functions of the first kind, and their argument is
$x=2V_{4}\,e^{-8\pi^{2}/g^{2}}|K|.$
If we sum these expressions in (18)-(19) over $\Delta n$ and evaluate the ratio on the right-hand side of (17), we recover the result (15) for the connected correlator $\langle{\rm Correlator}\rangle_{\rm conn}$. It is important to stress, that there are no infinities appearing in the final expression on the right hand side of (15) associated with an infinite volume. The integration over the instanton position $x_{0}$ is soaked up by the $(x_{i}-x_{0})$ and $(x_{i}-x_{0})$ arguments of the two fermion zero modes in the correlator $\langle\psi(x_{i})\bar{\psi}(x_{j})\rangle_{1}$ (or for the anti-instanton, in $\langle\psi(x_{i})\bar{\psi}(x_{j})\rangle_{\bar{1}}$) and no explicit factors of $V_{4}$ are left in (15). So, in the standard statistical mechanics-inspired ideal instanton gas approach, one keeps the volume finite: as already noted, the sums in (4) and (12) converge for any finite value of $x$, and the dependence on the spacetime volume cancels in the ratio appearing in (15).
The approach advocated in Refs. [3, 4] is different. These authors insist on taking the infinite volume limit before the infinite sum over $\Delta n$. Specifically, they define the correlator,
$\langle{\rm Correlator}\rangle_{\rm ACGT}\,=\,\lim_{N\to\infty}\,\lim_{V_{4}\to\infty}\frac{\sum_{\Delta n=-N}^{N}\sum_{n}\,\langle{\rm Correlator}\rangle_{n,\,\bar{n}=n-\Delta n}}{\sum_{\Delta n=-N}^{N}\sum_{n}\,Z_{n,\bar{n}=n-\Delta n}}$ (20)
imposing the $\lim_{V_{4}\to\infty}$ prior to extending the summation over $\Delta n$ to an infinite domain. In these settings, one makes use of the asymptotic values of the Bessel functions,
@@
so that all ratios of $I_{\Delta n\pm 1}(x)/I_{m}(x)$ and $I_{\Delta n}(x)/I_{m}(x)$ are universal and =1=1 in the $V_{4}\to\infty$ limit. Normalising the numerator and denominator expressions (18)-(19) by the same $I_{m}(x)$ and taking the limit $x\to\infty$, one obtains for the numerator of (20):
@@
The authors of Refs. [3, 4] take this cancellation to imply that physical observables, as defined by the order of limits in (20), are insensitive to the CP-violating parameter $\theta$. This assertion forms the core of their claim that CP violation is absent in QCD.
@@
    Since the framework of Refs. [3, 4] does not attempt to go beyond the ideal, non-interacting instanton gas and does not incorporate genuine multi-instanton contributions, it is particularly puzzling that the instanton correlators arising in their construction differ from the leading, single-instanton contributions on the right-hand side of (15).
@@
    The rationale in Refs. [3, 4] for taking the large-$V_{4}$ limit before summing over all topological sectors $\Delta n$ was to localise instantons well inside $V_{4}$, sufficiently far from the boundary at infinity. This was intended to ensure that, in every topological sector, the gauge fields approach the required pure-gauge boundary conditions.
@@
    In the presence of scalar field VEVs the gauge fields are massive and in practice one employs the so-called ‘constrained’ instantons [1] whose gauge fields decay as $A_{\mu}\sim e^{-m|x|}$ at large distances outside the instanton core. In this case instantons are always automatically localised within the spacetime volume $V_{4}$: the gauge fields vanish at infinity rather than approaching a pure-gauge configuration. Consequently, there is no valid justification for taking the infinite-volume limit prior to summing over distinct instanton-charge sectors.
@@
    Theta-dependence beyond QCD: Instanton contributions in 4-dimensional gauge theories always carry complex phases which are naturally organised as a Fourier series in $q=e^{2\pi i\tau}$, where $\tau=\frac{\theta}{2\pi}+\frac{4\pi i}{g^{2}}$. In the Seiberg–Witten $\mathcal{N}=2$ supersymmetric $SU(2)$ gauge theory [28], the low-energy effective action is encoded in the holomorphic prepotential $\mathcal{F}(a,\tau)$ where $a$ is the VEV of the $\mathcal{N}=1$ chiral superfield $\Phi$. For large $|a|\gg|\Lambda_{\scriptscriptstyle\mathcal{N}=2}|$, the theory is weakly coupled and one obtains an expansion
    $\mathcal{F}(a,\tau)$ $=$ $\mathcal{F}_{\mathrm{\rm pert}}(a,\tau)+\sum_{k\geq 1}\mathcal{F}_{k}(a)\,e^{2\pi ik\tau}$ (29)
    $=$ ${i\over 2\pi}a^{2}\,\log{2a^{2}\over e^{3}\Lambda_{\scriptscriptstyle\mathcal{N}=2}^{2}}\ -\ {i\over\pi}\sum_{k=1}^{\infty}F_{k}\,\left({\Lambda_{\scriptscriptstyle\mathcal{N}=2}\over a}\right)^{4k}a^{2}$ (30)
    where the summation is over all instanton numbers $k$. Multi-instanton coefficients $F_{k}$ were computed from first principles in Refs. [17, 16, 25] and were shown to agree precisely with the Seiberg–Witten predictions [28]. The complex phases $e^{2\pi ik\tau}$ of the multi-instanton contributions appearing in line (29) are essential for ensuring holomorphy and for reproducing the correct monodromy structure of the prepotential. The complex-valued dimensional transmutation scale $\Lambda_{\scriptscriptstyle\mathcal{N}=2}$ in (30) is given by the standard RG-invariant (in the supersymmetric setting) expression
    $\Lambda_{\scriptscriptstyle\mathcal{N}=2}^{b_{0}}\,=\,\mu^{b_{0}}\,e^{2\pi i\tau(\mu)}\,=\,\mu^{b_{0}}\,e^{-\frac{8\pi^{2}}{g^{2}(\mu)}+i\theta}\,,\qquad b_{0}=2N_{c}=4\,$, (31)
    where $\mu\in\mathbb{R}$ is the RG scale, which may be chosen as $\mu=|a|$. The dependence on the $\theta$-parameter in line (30) therefore remains explicit and is encoded in the phase of $\Lambda_{\scriptscriptstyle\mathcal{N}=2}$. This phase can be combined with the phase of the scalar VEV $a$ into a single physical CP-odd parameter,
    $\bar{\theta}\;=\;\theta-4\,\arg a\,$, (32)
    which is invariant under anomalous $U(1)_{R}$ transformations.
@@
    Turning now to the case of $\mathcal{N}=4$ SYM theory, we note that the factors of $e^{2\pi ik\tau}$ allow non-perturbative correlators to be organised into genuine modular or quasi-modular forms [6, 19] and [15]. They implement S-duality by providing the correct transformation behaviour under $SL(2,\mathbb{Z})$.
    We see that $\theta$-phases are not a fantom, instead they play an essential role in the non-perturbative QFT dynamics based on well-established exact results in supersymmetric YM models.
  6. 6.
    Strong CP from SUSY: To actually generate a $\bar{\theta}$-dependent vacuum energy, $E(\bar{\theta})$, aka the axion potential, we have to softly break supersymmetry. [^5] Otherwise the vacuum energy is trivially vanishing in any supersymmetric vacuum due to an exact cancellation between bosonic and fermionic contributions. Arguably, the simplest starting point is an $\mathcal{N}=1$ supersymmetric $SU(2)$ pure gauge theory. The theory has a mass gap and there are two (confining) supersymmetric vacua, $|{\rm vac}\rangle_{\pm}$ characterised by the corresponding values of the gluino condensate,
    $\frac{1}{16\pi^{2}}\,\langle{\rm tr}\lambda^{2}\rangle_{\pm}\,=\,\pm\,\Lambda_{\scriptscriptstyle\mathcal{N}=1}^{3}\,=\,\pm\,e^{i\frac{1}{2}\theta}\,|\Lambda_{\scriptscriptstyle\mathcal{N}=1}|^{3}.$ (33)
    where the complex-valued dimensional transmutation scale of the $\mathcal{N}=1$ theory is (cf. eq. (31)),
    $\Lambda_{\scriptscriptstyle\mathcal{N}=1}^{3}\,=\,\mu^{3}\,e^{\,\frac{1}{2}2\pi i\tau(\mu)}\,=\,\mu^{3}\,e^{-\frac{4\pi^{2}}{g^{2}(\mu)}\,+\,i\frac{1}{2}\theta}\,,\quad b_{0}=3N_{c}=6$ (34)
    Since $b_{0}=6$ in our model, the expression for $\Lambda_{\scriptscriptstyle\mathcal{N}=1}^{3}$ in (34) corresponds to a square root of a 1-instanton factor, or more precisely, to a contribution from a fractional instanton carrying instanton charge $1/2$.
    This reasoning was made precise in Refs. [12, 13] where the where the theory was compactified on $\mathbb{R}^{3}\times S^{1}$. The gauge field component along the compact direction acquires a vacuum expectation value $\langle A_{4}^{(3)}\rangle\,=\,v~\sim 1/R$, where $R$ is the radius of $S^{1}$. In the small radius regime, $R\ll 1/|\Lambda_{\scriptscriptstyle\mathcal{N}=1}|$, the theory is weakly-coupled, with $v/|\Lambda_{\scriptscriptstyle\mathcal{N}=1}|\gg 1$, and a semi-classical approximation is therefore justified. The semi-classical configurations of minimal Euclidean action on $\mathbb{R}^{3}\times S^{1}$ are the monopole-instantons which carry the instanton charge $1/2$. For gauge group $SU(2)$ there are two such configurations on $\mathbb{R}^{3}\times S^{1}$, constructed in Refs. [24, 23], commonly referred to as the BPS and KK monopoles. Their semi-classical contributions the gluino condensate were computed in [12],
    $\left\langle{\rm tr}\lambda^{2}\over 16\pi^{2}\right\rangle_{\scriptscriptstyle\rm BPS\,mono}\ =\ \left\langle{\rm tr}\lambda^{2}\over 16\pi^{2}\right\rangle_{\scriptscriptstyle\rm KK\,mono}\ =\ {\textstyle{1\over 2}}\,\mu^{3}\,e^{-{4\pi^{2}\over g^{2}(\mu)}\,+\,i\frac{1}{2}\theta}\,=\,{\textstyle{1\over 2}}\,e^{i\frac{1}{2}\theta}\,|\Lambda_{\scriptscriptstyle\mathcal{N}=1}|^{3}.$ (35)
    Summing the contributions of both monopole configurations reproduces the value of the gluino condensate in (33). The relative sign of the condensate in the two vacua is recovered by shifting the $\theta$-parameter in (35) by $2\pi$. Importantly, since the gluino condensate is a holomorphic $F$-term, its value, computed in the weakly coupled regime at small $R$, can be analytically continued to the decompactified, strongly coupled theory by taking the limit $R\to\infty$. In this limit, the result (33) remains valid [12].
    There is also an alternative and complementary semiclassical approach: the weak-coupling instanton calculation of Novikov, Shifman, Vainshtein, and Zakharov [26]. In this framework, the gluino condensate (33) is computed in a theory augmented by additional matter fields, whose presence renders the model weakly coupled and justifies a conventional constrained instanton analysis. Holomorphy of supersymmetric $F$-terms is then exploited to decouple the extra matter fields and to flow to the original pure $\mathcal{N}=1$ gauge theory, once again reproducing the expression on the right-hand side of (33), including its crucial $\theta$-dependence.
    We now break supersymmetry to $\mathcal{N}=0$ by introducing a mass term for the gluino,
    $\Delta\mathcal{L}_{\rm soft}=m_{\lambda}\,\lambda\lambda+\text{h.c.}$ (36)
    The soft mass term above involves a complex-valued mass parameter $m_{\lambda}$, and the ordering, $|m_{\lambda}|\ll\Lambda_{\scriptscriptstyle\mathcal{N}=1}$, ensures that this is a small deformation of the original theory. As a result, the vacuum energy $E$, which vanishes in the supersymmetric theory, is lifted by the soft breaking terms in (36), yielding:
    $\Delta E_{\pm}\,=\,-\left(m_{\lambda}\langle\lambda\lambda\rangle_{\pm}\,+\,m_{\lambda}^{*}\langle\bar{\lambda}\bar{\lambda}\rangle_{\pm}\right)\,=\,\mp\,\varepsilon\,\cos\left(\bar{\theta}/2\right)$ (37)
    to first order in $m_{\lambda}$. Here we defined $\varepsilon\,=\,32\pi^{2}|m_{\lambda}\Lambda_{\scriptscriptstyle\mathcal{N}=1}|$ and introduced the physical $\bar{\theta}$ phase, which is invariant under chiral $U(1)_{R}$ rotations, $\bar{\theta}\,=\,\theta+2\,{\rm arg}\,m_{\lambda}$.
    The vacuum energy density is obtained by taking the lower envelope of the two branches,
    $E(\bar{\theta})\,=\,{\rm min}\,\left(-\,\varepsilon\cos(\bar{\theta}/2)\,,\,+\,\varepsilon\cos(\bar{\theta}/2)\right)$ (38)
    As $\bar{\theta}$ is varied, a first-order phase transition occurs at $\bar{\theta}=\pi$, where the theory switches from one branch to the other.
    Figure 1: The figure on the left displays the two branches of the vacuum energy in (37). The physical (combined multi-branch) vacuum energy density $E(\bar{\theta})$ of the $SU(2)$ YM theory of (38) is shown on the right. $E(\bar{\theta})$ is a periodic function with a period $2\pi$ and cusps at $\bar{\theta}=\pi$ (mod $2\pi$).
    The resulting multi-branch structure $E(\bar{\theta})$, or equivalently of the axion potential, (38) is plottted in the right-hand-side panel of Fig. 1. Our simple $SU(2)$ Yang–Mills model thus provides a concrete realisation of Witten’s general multi-branch periodic potential, originally motivated in the large-$N_{c}$ limit by Witten, Di Vecchia, and Veneziano in Refs. [29, 14].
  7. 7.
    Strong CP in the softly broken Seiberg-Witten theory: It is also well-known that the same non-trivial (multi-brach) periodic structure of $E(\bar{\theta})$ follows from the $\mathcal{N}=2$ super–Yang–Mills theory broken to $\mathcal{N}=0$ either by introducing soft gluino mass terms [22] or by turning on an anomaly mediated supersymmetry breaking [10] The Coulomb branch of the Seiberg-Witten model is parametrised by the flat direction
    $u\equiv\langle{\rm tr}\Phi^{2}\rangle,$ (39)
    where $\Phi$ is the $\mathcal{N}=1$ chiral superfield. SUSY breaking $\mathcal{N}=2$ to $\mathcal{N}=1$ is achieved by adding the mass-term for $\Phi$,
    $W_{\mathcal{N}=1}=m\,{\rm tr}\Phi^{2},$ (40)
    in which case the flat direction is lifted producing two isolated $\mathcal{N}=1$ vacua, located near the Seiberg–Witten singularities
    $u_{\pm}\,=\,\langle{\rm tr}\Phi^{2}\rangle_{\pm}\,=\,\pm\,\Lambda_{\scriptscriptstyle\mathcal{N}=2}^{2}$ (41)
    Next step is to break $\mathcal{N}=1$ to $\mathcal{N}=0$ by soft gluino masses, as in (36). The vacuum energy of the resulting pure $SU(2)$ gauge theory is given by [22],
    $E(\bar{\theta})\,=\,{\rm min}\,\left(-\,\varepsilon\cos(\bar{\theta}/2)\,,\,+\,\varepsilon\cos(\bar{\theta}/2)\right),$ (42)
    where the energy-density scale $\varepsilon$ is now $\varepsilon\,=\,32\pi^{2}|m\,m_{\lambda}\Lambda_{\scriptscriptstyle\mathcal{N}=2}|$ and the invariant CP-odd phase is $\bar{\theta}\,=\,\theta+2\,({\rm arg}\,m_{\lambda}\,+\,{\rm arg}\,m)$. The functional form of $E(\bar{\theta})$ in (42) is identical to that obtained above in (38) and plotted in Fig. 1. The result is derived in the parametric regime $|m_{\lambda}|\ll|m|\ll|\Lambda_{\scriptscriptstyle\mathcal{N}=2}|$, where the soft breaking scales may be treated as small deformations of the underlying Seiberg–Witten theory.
    This construction also provides a natural explanation for the multi-branch structure of the vacuum energy $E(\bar{\theta})$. The weakly coupled degrees of freedom in the neighbourhood of the two vacua $u_{\pm}$ are mutually non-local. The branch $\varepsilon\cos(\bar{\theta}/2)$ is associated with the $u_{+}$ vacuum, in which the $(1,0)$ monopole condenses, whereas the second branch, $\varepsilon\cos(\bar{\theta}/2)$, corresponds to the $u_{-}$ vacuum, where the condensing field is the $(1,1)$ dyon. For $\pi<\bar{\theta}<\pi$, the appropriate weakly coupled degrees of freedom are monopoles, while in the interval $\pi<\bar{\theta}<3\pi$ the relevant description involves dyons.
    In a subsequent study, Ref. [10] analysed the $\theta$-dependence of $\mathcal{N}=2$ $SU(2)$ QCD with $N_{f}$ fundamental flavours, explicitly broken to $\mathcal{N}=1$ and coupled to anomaly-mediated supersymmetry breaking. For $N_{f}=0$, this analysis reproduces the potential shown in Fig. 1, and generalises it to $N_{f}=1,2,3$, revealing cusps and associated phase transitions in $E(\bar{\theta})$ at fractional values of $\bar{\theta}$.
  8. 8.
    Conclusion -- when instantons are and are not enough: Instanton-based semiclassical methods are reliable only in weakly coupled theories. QCD in normal conditions does not satisfy this criterion. Consequently, the use of instantons to determine the $\theta$-dependence of physical observables requires modifying the theoretical setting so as to access a weakly coupled phase. One may then apply instanton-based techniques and subsequently assess whether the resulting insights can be meaningfully extrapolated to the original strongly coupled theory. A phenomenologically relevant example is provided by QCD at high temperature, $T\gg\Lambda$. In this regime the theory becomes weakly coupled, the instanton formalism [20] is well justified, and the vacuum energy $E(\bar{\theta};T)_{\rm QCD}$ can be reliably computed using the dilute instanton gas approximation (DIGA).
    Another class of controlled modifications involves introducing scalar fields with vacuum expectation values much larger than the strong-coupling scale, $\langle\phi\rangle\gg\Lambda$. In such settings (which include applications to the electroweak theory and to supersymmetric models) one employs constrained instanton configurations [1]. These can be used to compute, for example, the Seiberg–Witten prepotential [17, 16, 25] or the gluino condensate [26]. The resulting expressions can then be analytically continued to the strongly coupled regime, as discussed above, thereby recovering the $\theta$-dependence of the gauge theory.
    One can also consider compactifying compactifying spacetime on $\mathbb{R}^{3}\times S^{1}$, which allows the use of fractionally charged instanton–monopole configurations [12]. Other more recent applications of semiclassical techniques to the computation of the $\theta$-dependence and axion potentials based on small instanton-like configurations in gauge theories, can be found in Refs. [2, 11, 18].
    In all of the scenarios considered above, the $\theta$-dependence is non-trivial.
In this note, I have first focused on instanton-based derivations of the $\theta$-dependence in theoretical settings where the gauge dynamics is weakly coupled, and subsequently on examples involving softly broken supersymmetric theories at strong coupling. For a selection of complementary formulations and alternative approaches, the reader is referred to the ongoing discussion in the recent literature, including Refs. [7, 5, 27].